OpenStudy (jjuden):

Find the volume of the box v=lwh L=4a-3 W=a+4 H=2a-2

OpenStudy (anonymous):

Volume = length *width * height

OpenStudy (jjuden):

i did thats but it wasnt right

OpenStudy (anonymous):

\[8a ^{3}-8a ^{2}\]

OpenStudy (the_fizicx99):

Did you try, V = (4a-3)(a+4)(2a-2) ?

OpenStudy (jjuden):

yes i tried thats what else would i try

OpenStudy (jjuden):

@felavin how did u do that

OpenStudy (anonymous):

this is simple, you just multimple LxWxH

OpenStudy (jjuden):

its not simple :(

OpenStudy (the_fizicx99):

V = (4a-3)(a+4)(2a-2) V = (4a^2 + 16a -3a -12)(2a-2) V = 8a^3 - 8a^2 + 32a^2 -32a -6a^2 + 6a -24a + 24 V = 8^3 + 18a^2 -50a + 24

OpenStudy (jjuden):

ohhhhhhhh ok i get it thank you for explaining

OpenStudy (the_fizicx99):

Yw :>